cos^2ydx+(1+e^-x)sinydy=0
1
Expert's answer
2020-07-07T21:16:30-0400

Diffrential equation is

"\cos^2(y)dx+(1+e^{-x})\sin(y)dy=0"

Now, divide the above equation by "\cos^2(y)(1+e^{-x})" ,thus

"\frac{dx}{1+e^{-x}}+\frac{\sin(y)}{\cos^2(y)}dy=0\\
\implies \frac{e^x}{1+e^x}dx+\frac{\sin(y)}{\cos^2(y)}dy=0\\
\implies \int\frac{e^x}{1+e^x}dx+\int\frac{\sin(y)}{\cos^2(y)}dy=\text{constant}\\
\implies \ln(1+e^x)-\int\frac{1}{u^2}du=\text{constant, where $u=\cos(y)$}\\
\implies \textcolor{blue}{\ln(1+e^x)+\frac{1}{\cos(y)}}=\text{\textcolor{blue}{constant}}"